OpenStudy (anonymous):

how do u find t in d= vt+1/2 at^2?

OpenStudy (tkhunny):

It's a quadratic equation. Do you recall complaining about solving SO MANY of those?! Now's your chance to prove that you get it!

OpenStudy (kc_kennylau):

You plug all the values in and solve the quadratic equation

OpenStudy (anonymous):

ii dont get it... :'( coz theres 2 ts in it so u cnat rearrange the formula to find t....

OpenStudy (tkhunny):

You didn't listen to me. Do you remember the quadratic formula? You can do it!!

OpenStudy (anonymous):

i think ive rearraged the formula tp 2(t^2 + 32t - 70)=0

OpenStudy (tkhunny):

If you were given d, v, and a, so that you have this result, then what is now preventing you from solving it?

OpenStudy (anonymous):

i cant solve it though, argh im so dumb

OpenStudy (tkhunny):

What class are you in?

OpenStudy (kc_kennylau):

Do you know how to solve quadratic equations? There are three ways of solving one.

OpenStudy (tkhunny):

...or more.

OpenStudy (anonymous):

i do but idk what multiples of 70 can add to 32

OpenStudy (tkhunny):

Why are you trying to factor? Don't do that.

OpenStudy (tkhunny):

What were you given for d, v, and a?

OpenStudy (kc_kennylau):

If no multiples of 70 can add to 32, you can use the second or the third way.

OpenStudy (kc_kennylau):

The second way is completing the square, the third way is quadratic formula

OpenStudy (anonymous):

d=120 v=32 a=9.8

OpenStudy (kc_kennylau):

d=vt+1/2 at^2 120 = 32t + 1/2 9.8t^2 120 = 32t + 4.9t^2

OpenStudy (kc_kennylau):

4.9t^2+32t-120=0 49t^2+320t-1200=0

OpenStudy (kc_kennylau):

How the hell did you get 2(t^2 + 32t - 70)=0

OpenStudy (anonymous):

lol i think im too tired.... gonna take a rest

OpenStudy (anonymous):

You should listen to tkhunny, he asked you a couple of times, do you recall the quadratic equation... but you seem to keep ignoring him, to solve t you need to use it.